### v2

# Aggregation in dilute aqueous solutions of hydroxypropyl cellulose with salt ions

Aggregation in dilute aqueous solutions of hydroxypropyl cellulose with salt ions
Valery I. Kovalchuk
Taras Shevchenko National University of Kyiv, Faculty of Physics
(64/13, Volodymyrska Str., Kyiv 01601, Ukraine)
Keywords: static and dynamic light scattering, hydroxypropyl cellulose, salt ions, Cahn-Hilliard equation.
1. Introduction
The study of the solutions of cellulose derivatives is based on promising technologies in various fields, including the food industry, construction, hydrocarbon production, aerospace materials, medicine and pharmaceuticals [1, 2, 3]. A characteristic feature of many water-soluble cellulose ethers is the thermoreversible volume transition [4]. Its essence consists in the formation of a polymer gel network when the temperature increases, but the system returns to the state of isotropic solution when cooled down. The threshold temperature of this transition is called the lower critical solution temperature (LCST) which depends on a number of factors such as polymer concentration, type and degree of substitution [5, 6, 7], pH value [8, 9, 10, 11], and the presence of electrolytes in solution [12, 13, 14, 15, 16].
A number of works are devoted to the study of the thermosensitive behavior of cellulose derivatives aqueous solutions with salt ions [17, 12, 13, 14, 18, 19, 20, 21, 22]. The main methods that were used: turbidimetry [14, 18, 19, 22], dynamic light scattering (DLS) [17, 19, 21, 22], microcalorimetry [12, 13, 14, 18], and viscometry [13, 14, 18, 20, 21, 22]. In the articles listed above the volume phase transition mechanisms were mainly studied depending on the type of salt additives and their concentration. In particular, it was shown that the thermosensitive behavior of polymer solution is based on competition for water molecules between polymer chains and salt ions, which leads to the hydrophobic aggregates formation [12, 13, 14, 15].
It should be noted that ions can act as aggregation initiators not only in solutions of cellulose derivatives, but also in other systems containing, for example, peptides [23], pseudo-polypeptoids [24], lignin [25], lipid nanoparticles [26], latex particles [27, 28], and silica particles [29]. In general, the study of aggregation in polymer solutions is of great practical importance for modern technologies, such as catalysts with controlled activity and film nanocomposite materials [30, 31, 32], color-based sensors that respond to pH and heavy metal ions [33], energy-saving smart windows [34, 35], and drug delivery systems [8, 9].
Hydroxypropyl cellulose (HPC) as an object of research was chosen due to the attractiveness of its physical and chemical properties. In addition to the well-known unique properties of a green polymer, such as availability, cheapness, and biocompatibility, HPC has a low LCST value of $(40\div 45)^{\circ}$ [34, 35, 8, 9, 36, 37, 7]. Such a temperature makes HPC a good choice for both biomedical applications and promising smart technologies [10]. In our recent work [22], it was found that the addition of salts (chlorides of Group I alkali metal ions) to dilute aqueous solutions of HPC leads to polymer aggregation at temperatures above LCST. As a continuation of [22], in this work these same systems are studied using static light scattering. The obtained light backscattering data were analyzed within the framework of the tangent plane approximation and Mie scattering theory. It is also shown that the hydrophobic association of polymer molecules in the presence of electrolytes in solution can be described within the framework of the Cahn-Hilliard model.
2. Materials and Methods
2.1. Materials
The cellulose derivative, hydroxypropyl cellulose (HPC), was purchased from Alfa Aesar company. The manufacturer’s specification indicates that HPC has an average degree of substitution of 75.7% and a molecular weight of 100,000. The viscosity range was reported by the manufacturer to be 112 cPs at $25^{\circ}$ for a 5 wtaqueous solution.
Analytical grade salts (Li, Na, K, Rb, Cs chlorides) were purchased from Sigma-Aldrich company.
2.2. Specimen preparation
The initial aqueous HPC solution with a concentration of 2 wtwas prepared by dissolving the necessary amount of polymer in deionized water under continuous stirring for 4 hours at a temperature of $60^{\circ}$ to ensure the complete dissolution of the polymer. As a result, a homogeneous and transparent solution was obtained.
This initial solution was divided to prepare six specimens. The salts were introduced into five of them, and the specimens were mixed until the complete salt dissolution. The molar concentrations of the salt in the specimens were identical and equal to that of the physiological solution (154 mmol/l). By diluting 2% solutions with water to a ratio of 1:10, specimens with a polymer concentration of 0.2 wtand a salt concentration of 15.4 mmol/l were fabricated.
2.3. Dynamic light scattering
Particle sizes (hydrodynamic diameters) were determined using a Zetasizer Nano ZS (Malvern, UK) instrument at $173^{\circ}$ backscatter geometry. Before measurement, all specimens were dust-removed through a 0.45 um filter (Minisart NML). For each temperature point, the particle size distributions in a given specimen were measured three times.
2.4. Static light scattering
Fig. 1 shows the schematic of the measuring chamber of the experimental setup used for studies of polymer solutions by static light scattering [38, 39].
Figure 1: Schematic of the measuring chamber of the experimental setup.
As light source, we used a GNL-5013PGC LED with a known spectral characteristic. It was powered using a micropower current stabilizer made on an LP2951 chip. Light with a wavelength of 525 nm was fed through fiber optic cable 1 into thermostatic chamber 2 filled with the polymer solution. The light beam reflected from mirror 3 returned through optical fiber cable 4 to a digital photosensor connected to an ATmega328P microcontroller. As a result, the total intensity
$J_{\Sigma}=J_{\text{T}}+J_{\text{R}}$ (1)
was measured, where $J_{\text{T}}$ is the intensity of the light flux attenuated because of its passage through the specimen, and $J_{\text{R}}$ is the intensity of the light flux scattered by the specimen at an angle of $180^{\circ}$. When measuring $J_{\text{R}}$, mirror 3 was screened with curtain made of light-absorbing material (black anodized aluminum). Thus, in order to determine $J_{\text{T}}$ and $J_{\text{R}}$ for a specific specimen, it was necessary to perform two measurements: in the presence and in the absence of screen.
The solution temperature was measured by means of a digital temperature sensor LMT01LPG 5 and another an ATmega328P microcontroller. The illumination and temperature values were synchronously read out using a microcontrollers and transferred to USB ports of a personal computer using the RS-232 protocol. Accumulation of data, their further processing and visualization were implemented with the help of a Delphi code. The illuminance measurement error did not exceed 0.1 lux, and the temperature measurement error was ±$\pm 0.1^{\circ}$. The chamber with the polymer solution was placed in light-proof box and connected with the help of nozzles 6 and 7 to a circulation thermostat Julabo ME-6. Structurally, elements 3 and 5 together with the output (input) of optical fiber cables 1 and 4 were arranged in the form of a probe that was immersed into the examined solution.
3. Results and Their Discussion
3.1. Static and dynamic light scattering
When falling on the surface of the examined solution (see Fig. 1), the primary beam of light formed two beams, transmitted and reflected. The corresponding intensities $J_{\text{T}}$ and $J_{\text{R}}$ of these beams were measured at a set of temperatures within an interval from 30 to $65^{\circ}$. The specimen heating rate was $1.1^{\circ}$/min. The obtained experimental data were calibrated in each specimen to a transparency level of 100%. As such the corresponding value of $J_{\text{T}}$ at a temperature of $30^{\circ}$ was selected.
Figure 2 exhibits the measurement results obtained for the temperature dependences of the relative intensities of the backscattered light beam for all six specimens. The same figure also shows the average size of polymer particles (hydrodynamic diameter, $D$) determined by the dynamic light scattering method (points) [22].
Figure 2: Temperature dependences of the reflected light intensity $J_{\text{R}}$ (left vertical scale, $\circ$) for specimens: (a) – ion-free, (b) – Li, (c) – Na, (d) – K, (e) – Rb, (f) – Cs. The hydrodynamic diameter values of the polymer clusters $D$ are also shown here (right vertical scale, $\bullet$).
3.2. Light Backscatter Analysis
Fig. 2 shows that the experimental dependences $J_{\text{R}}(T)$ for the solutions with salt ions have a non-monotonic character: as the temperature increases, the $J_{\text{R}}$-values first grow, then reach a maximum, and afterward decrease. This behavior can be explained as follows.
The solution surface on which the light falls is not perfectly smooth – it contains polymer inclusions of various sizes, therefore it can be considered as a statistically rough surface. The studies of light scattering by such surfaces have a long (for more than a hundred years) story [40]. It is known that the solution of this problem is reduced to the analysis of a wave equation of a certain type. Instead, in this paper, we will confine ourselves to a qualitative consideration proceeding from assumptions concerning the type of backward scattering, which forms the output beam with the intensity $J_{\text{R}}$.
Let the solution surface look like a plane on average. Let’s draw the $z$-axis perpendicular to this plane. Surface roughness $z=\xi(r)$ is described by two statistical parameters [41]: parameter $\delta$ (root-mean-square height of surface deviation from the plane $z=0$) and parameter $\alpha$ (lateral correlation length of roughness, describing the average lateral distance between the peak and valley of the surface profile).
We assume that the incident light propagates along the $z$-axis and that the condition
$(kR_{\text{c}})^{1/3}\gg{1}$ (2)
is satisfied. Here $k$ is the wave vector modulus of the photon, $R_{\text{c}}$ is the radius of surface curvature, which can be estimated as $R_{\text{c}}\simeq 2\sqrt{3}\alpha^{2}/\delta$ [42].
Inequality (2) is a condition for the applicability of the tangent plane method, developed in [43] based on the theory of wave scattering by a statistically rough surface [40]. The main points and conclusions of this method as applied to backscattering of light are as follows.
When considering methods for assessing the reflective properties of light by a rough surface, the entire reflected signal is divided into a coherent component and an incoherent one. The coherent component is associated with specular reflection from surface areas and is determined by the average value of the field strength in the scattered wave. According to [40], the back reflection coefficient of the coherent component for a surface with a normal distribution of heights is equal to
$\rho=J_{\text{R}}/J_{0}=\exp(-2k^{2}\delta^{2}),$ (3)
where $J_{0}$ is the intensity of incident light beam. From (3) it follows that the value of $J_{\text{R}}$ decrease exponentially as the characteristic height of irregularities increases.
In [40, 44] it was shown that the cross section of backscattering of light by a three-dimensional statistically rough surface is
$\Sigma\sim{Na^{2}},$ (4)
where $N$ is the number of scattering centers (points of specular reflection, which correspond to the peaks and valleys of the surface $z=\xi(r)$), and the value $a$ is the geometric mean of the main radii of curvature $a_{1}$ and $a_{2}$ at the points of specular reflection
$a^{2}=\langle{a_{1}a_{2}}\rangle.$ (5)
Formula (4) has a simple geometric meaning: the scattering cross section by a rough surface in the tangent plane approximation (2) coincides with the scattering cross section by $N$ identical balls with radius $a$. Consequently, the problem of light scattering by a statistically rough surface can be reduced to the equivalent problem of scattering by a system of spherical particles. Since the amount of polymer in the solution is a constant parameter, the number of such particles equals
$N\sim 1/a^{3},$ (6)
and, therefore,
$\Sigma\sim 1/a.$ (7)
Note that the result (7) can also be obtained in the framework of Mie scattering theory [45]. The Mie backscattering cross section for a single particle has the form [46]
$\sigma(a)=\frac{1}{a^{2}k^{2}}\Bigl{|}\sum_{n=1}^{\infty}(-1)^{n}(2n+1)(A_{n}(a,k)-B_{n}(a,k))\Bigr{|}^{2},$ (8)
where $A_{n}$, $B_{n}$ are the Mie coefficients[47].
Figure 3 shows the computing results of $\sigma(a)$ that were performed using the [48, 49] method for light with a wavelength of 525 nm and a refractive index value for 0.2% water solution of HPC taken from [50].
Figure 3: Mie backscattering cross section for a single spherical aggregate of hydroxypropyl cellulose with radius $a$ (set of black dots).
This figure shows that the $\sigma(a)$ dependence has the scaling behavior (dashed line)
$\sigma(a)\sim{a^{s}}$ (9)
with the power exponent $s=2$. Hence, the total cross section of Mie backscattering for $N$ identical spherical particles, taking into account (6), is
$\Sigma_{\text{Mie}}\sim{Na^{2}}=1/a,$ (10)
which coincides with the formula (7). It follows from formula (10) that
$J_{\text{R}}\sim{1/a},$ (11)
that is, the intensity of backscattered light increases as the particle size decreases. The same conclusion can be reached by comparing the experimental values of $J_{\text{R}}$ and $D$ in Figure 2a with the values of the same quantities in any of the remaining Figures 2b-f.
As mentioned above, the behavior of $J_{\text{R}}(T)$ in Fig. 2b-f has a non-monotonic character: as the temperature increases, the $J_{\text{R}}$-values first grow, then reach a maximum at $T{\kern 1.0pt}^{\prime}\simeq(50\div 55)^{\circ}$, and afterward decrease. The sizes of the polymer clusters (see the same figures) have a similar temperature behavior. There are two explanations for the decrease of cluster sizes at $T>T{\kern 1.0pt}^{\prime}$. The first of them consists in that the clusters collapse at $T>T{\kern 1.0pt}^{\prime}$ and form numerous small fragments. But in this case, as follows from Eq.(11), instead of the reduction of the intensity $J_{\text{R}}$, we would observe its growth, which contradicts the experiment. The second explanation is that large polymer aggregates disappear from the solution at $T>T{\kern 1.0pt}^{\prime}$ owing to their sedimentation, which is confirmed by the experiment: whenever the measurements were concluded, some polymer sediment was found at the bottom of the cell with the solution. This sediment did not dissolve well in a hot solution, but it dissolved easily in cold water.
Note that the viscosity of the 0.2-wtaqueous solutions of HPC at $(50\div 60)^{\circ}$ exceeds the viscosity of water by only 5–7% [21]. Therefore, the sedimentation of polymer clusters occurred quite quickly: from Fig. 2b-f, it follows that $J_{\text{R}}$ decreased by a factor of two during about 10 min.
Figure 2 also demonstrates that the sizes $D$ of clusters in the ion-free specimen are several times smaller than in the specimen with ions. By order of magnitude, they are equal to the wavelength of visible light. In Fig. 2a, the intensity of reflected light is larger than that in Fig. 2b-f, as it should be according to formula (11). The polymer clusters in the ion-free solution were in a suspended state, and practically no sediment was observed in the cell after the measurements.
3.3. Computer simulation of HPC aggregation: the role of hydrophobic interactions
The hydrophobic association of the polymer in the vicinity of LCST can be investigated within the framework of the Cahn-Hilliard model developed from the free energy density functional method.
The Ginzburg-Landau type functional for the total energy of the polymer-solvent system has the form [51]
[…]
where $\phi$ is the order parameter meaning the polymer concentration in solution, $F(\phi)$ is the free energy density. The second term in curly brackets under the integral sign in (12) describes the contribution of spatial correlation effects to the free energy with a gradient coefficient [52]
$\kappa(\phi)=\frac{a^{2}}{36\phi(1-\phi)},$ (13)
where $a$ is the size of the polymer chain segments, equal to 1 for the Flory model [53].
Let’s introduce an additional parameter $\gamma$ into (12):
[…]
the physical meaning of which is as follows.
In the works of Cahn and Hilliard [54, 55] it was shown that the interfacial surface tension coefficient
[…]
Macromolecules of cellulose derivatives are amphiphilic, i.e. they contain both nonpolar hydrophobic and polar hydrophilic sites in their structure. The latter compete with salt ions for water molecules in solutions with electrolytes [19, 12, 13, 14]. As a result, the number of hydrophilic water-polymer bonds decreases and the interfacial energy density, respectively, increases. In other words, the enhancement of hydrophobic interactions at the polymer-solvent interface leads to an increase in $\sigma_{\text{i}}$. Thus, by changing the value of the parameter $\gamma$ in the formula (14), the hydrophobicity of the polymer can be simulated.
The Cahn-Hilliard equation describing the evolution of the polymer concentration at a point in space $\vec{r}$ at time $t$ has the form [51]
$\frac{\partial\phi}{\partial{t}}=\nabla\Bigl{\{}M\nabla\frac{\delta{U[\phi]}}{\delta\phi}\Bigr{\}}+\xi,$ (16)
where $M$ is the mobility, $\xi=\xi(\vec{r},t)$ is a stochastic function (thermal noise) satisfying the fluctuation-dissipative theorem [56]. To solve equation (16) with the functional (14), we use the Flory-Huggins free energy [53, 57] in the form
$F(\phi)=N^{-1}\phi\ln\phi+(1-\phi)\ln(1-\phi)+\chi\phi(1-\phi),$ (17)
where $N$ is the degree of polymerization of the molecular chain, $\chi$ is the Flory-Huggins parameter describing pairwise interactions between monomers. Then the variational derivative of the total energy in (16) is equal to
$\frac{\delta{U[\phi]}}{\delta\phi}=f(\phi)-\gamma\bigl{(}\lambda(\phi)|\nabla\phi|^{2}+2\kappa(\phi)\Delta\phi\bigr{)},$ (18)
where
$f(\phi)=N^{-1}(1+\ln\phi)-\ln(1-\phi)-\chi(2\phi-1)-1,$ (19)
$\lambda(\phi)=\frac{2\phi-1}{36\phi^{2}(1-\phi)^{2}}.$ (20)
We consider that the mobility of polymer molecules depends on their concentration
$M=M_{0}\phi(1-\phi).$ (21)
Finally, the Cahn-Hilliard equation takes the form
$\frac{\partial\phi}{\partial{t}}=M_{0}\nabla\Bigl{\{}\phi(1-\phi)\nabla\bigl{(}f(\phi)-\gamma(\lambda(\phi)|\nabla\phi|^{2}+2\kappa(\phi)\Delta\phi)\bigr{)}\Bigr{\}}+\xi.$ (22)
This is a fourth-order nonlinear parabolic partial differential equation with a stochastic term. To enforce the mass conservation law and descending of total energy with time, equation (22) must be equipped with the homogeneous Neumann boundary conditions [58].
Without loss of generality, let us investigate solutions $\phi(x,t)$ of the one-dimensional equation (22). Introducing a spatio-temporal grid for $x$ and $t$ with periodic boundary conditions, we use for solving of (22) the semi-implicit difference scheme [51] with the parameters of the polymer-solvent system under study taken from [39, 59]. The initial phase was homogeneous, and the corresponding polymer concentration for a 0.2 wtsolution was $\phi_{\text{in}}=2\cdot 10^{-3}\rho$, where $\rho=1.3\,\,\text{g/cm}^{3}$ is the density of hydroxypropyl cellulose [60]. The phase structure growth was simulated on a lattice with $L=128$.
In Figure 4, as an example, the result of numerical simulation is demonstrated for $\gamma=0.04$, where it is seen how the reduced concentration $\overline{\phi}=\phi/\phi_{\text{in}}$ changes depending on the dimensionless distance $x$ and time $t$. It can be seen how an unstable regime arises on the basis of a random distribution $(t=0.001)$, which leads to the primary structure of a new phase $(t=0.2)$, and then to its further coarsening and final formation $(t=1-5)$.
Figure 4: Distribution of the reduced concentration $\overline{\phi}(x,t)$ on a one-dimensional lattice during the spinodal decomposition in the hydroxypropyl cellulose-water system. The value of $\gamma$ is 0.04.
One of the features of the asymptotic behavior of the system under study at large times is dynamical scaling, which is determined by the characteristic length $L(t)$ of the individual ordered regions of the phase structure [61]. Choosing the average radius of cluster $R(t)$ as characteristic length at the level $\overline{\phi}=1$, we find the scaling exponent for the growth law $R(t)\sim{t^{\delta}}$. Analysis of the time dependencies $R(t)$ in their longest linear sections in Figure 5 gives the value $\delta=0.324\pm{0.026}$. This value is close to the exponent $\delta=1/3$ which is typical for the laws of growth in systems with a conserved scalar order parameter [62, 63].
Figure 5: Dependencies of the average cluster radius over time for the set of $\gamma$ values: 0.01 (1), 0.02 (2), 0.03 (3), 0.04 (4), 0.05 (5), 0.06 (6). Averaging was performed for 100 runs of the program simulating the growth of the phase structure.
Figure 6: Average cluster radius $R$ as a function of parameter $\gamma$. Simulation time is $t=5$.
Fig. 6 shows the dependence of average cluster radius $R$ on the parameter $\gamma$. It follows from Fig. 6 that the strengthening of hydrophobic interactions at the interfacial boundary (increase in the parameter $\gamma$) leads to an increase in the size of polymer aggregates.
4. Conclusions
Using the methods of static and dynamic light scattering, dilute aqueous solutions of HPC (0.2 wt%) with the admixtures of Group I alkali metal chlorides (15.4 mmol/l) are studied. It is found that, at temperatures higher than the LCST, there arises inverse scattering of light via the light reflection from the supramolecular structure of the researched solutions. In the solutions with ions, there emerge large macromolecular associations (clusters) whose size is several times larger than the wavelength of visible light. The temperature dependence of the intensity $J_{\text{R}}$ of backward scattering is non-monotonic for specimens with salts; namely, as the temperature grows, the values of $J_{\text{R}}$ increase, reach a maximum, and finally decrease. The hydrodynamic diameter of clusters is demonstrated the same behavior in corresponding solutions with ions, correlating with the $J_{\text{R}}$-values. There is a relation between the values of $J_{\text{R}}$ and the sizes of polymer clusters, which was found in the framework of the tangent plane approximation and Mie scattering theory. In particular, it is shown that the reduction of $J_{\text{R}}(T)$ in an interval of $(55\div 65)^{\circ}$ occurs due to the sedimentation of clusters rather than their decay, which was confirmed experimentally.
It is shown that in the vicinity of LCST the hydrophobic association of polymer molecules in solutions with salt ions can be described in the framework of the Cahn-Hilliard model. It follows from the results of computer simulation that the increase in the interfacial energy density at the polymer-solvent interface (strengthening of hydrophobic interactions) leads to an increase in the size of polymer aggregates.
This study may provide useful information for the development of polymer films with a given structure and characteristics, as well as new drug delivery systems.